company received a fixed asset as a donation and the market value of this fixed asset was$ 10,000. The depreciation rate of this asset is 20%pa. Asset was immediately put into use. The company sold this car after 2 years for the price of $ 5,000. What is the result of P/L section"profit(loss) on sale " in this comay inthe year of sale?
Answer #1

SOLUTION

given data

the value of this fixed asset was$ 10,000.

The depreciation rate of this asset is 20%pa.

then the company sold this car after 2 years for the price of $ 5,000.

When company receive asset on donation cost will be fair value  

cost of asset = 10000

depreciation for 2 years=10000*20%*2 4000

carrying value of asset=10000-4000 6000

sales price= 5000

loss on sale= 5000-6000 -1000

THANK YOU